OpenStudy (anonymous):

f(x)=3x^2+2x write the equation that shift the graph of this curve two units vertically and 3 units to the right simultaneosly

OpenStudy (solomonzelman):

vertical shift, lets take \[f(x)=2x\]it will be\[f(x)=2x+a,~~~a~~units~~up\]\[f(x)=2x-a,~~~a~~units~~down\]\[f(x)=2(x+a),~~~a~~units~~Left\]\[f(x)=2(x-a),~~~a~~units~~Right\]

OpenStudy (solomonzelman):

in your case it will be a little different because it's cubic, let me think....

OpenStudy (solomonzelman):

quadratic not cubic....

OpenStudy (solomonzelman):

I think I was wrong about horizontal shift. http://www.purplemath.com/modules/fcntrans.htm it is added or subtracted from f(x) like f(x+a)

OpenStudy (solomonzelman):

\(\color{black}{f(x)~=~3x^2~+~2x}\) \(\color{red}{~+2~~~~~~~Vertical ~~~shift~~2~~~units~~~up.}\)

OpenStudy (solomonzelman):

and\[f(x-3 )=3x^2+2x+2\]

OpenStudy (solomonzelman):

x-3 is a shfit to the right 3 units comparing to f(x) (without 3 units to the right)

OpenStudy (solomonzelman):

\(\huge\color{blue}{f~(x~\color{red}{-~3})~=~3x^2~+~2x~\color{red}{+~2})}\)

OpenStudy (anonymous):

thanks u r so helpful!